Outdoor air temperature is 0 ° C and relative humidity 90%. The air is blown into the house
and heated to 20 ° C. What is the new relative humidity?
1
Expert's answer
2019-03-14T17:32:55-0400

Solution:

Relative humidity is the ratio of vapor density to saturation vapor density.

Relative humidity of the air at "0^\circ C:"


"\phi_1=\frac{\rho} {\rho_{01}}100\%"

Relative humidity of the air at "20^\circ C:"

"\phi_2=\frac{\rho} {\rho_{02}}100\%"

"\rho_{01}=4.84 g/m^3,\: \rho_{02}=17.2 g/m^3" are saturation vapor densities at "0^\circ C\:and\:20^\circ C" respectively.

From first equation:

"\rho=\frac{\phi_1 \rho_{01}} {100\%}"

Therefore:

"\phi_2=\frac{\phi_1 \rho_{01}100\%} {\rho_{02}100\%}=\frac{\phi_1 \rho_{01}} {\rho_{02}}=\frac{90\%\cdot 4.84} {17.2}=25\%"

Answer:

"\phi_2=25\%"